OpenStudy (anonymous):

please solve 3(n+1)/7 is equal to or greater than n+4/5

OpenStudy (anonymous):

\[\frac{{3}(n+1)}{7}\geq\frac{(n+4)}{5}\] Playing with the equation editor. Still there?

OpenStudy (anonymous):

\[n \le -7\div5\]

OpenStudy (anonymous):

Hmm. I'm getting a different answer. Check my work: \[\frac{5}{1}*\frac{7}{1}*\frac{{3}(n+1)}{7}\geq\frac{5}{1}*\frac{7}{1}*\frac{(n+4)}{5} \]

OpenStudy (anonymous):

\[5(3n+3)\geq{7n+28}\Rightarrow15n+15\geq{7n+28}\]

OpenStudy (anonymous):

\[8n\geq13\Rightarrow{{n}{\geq}\frac{13}{8}}\]